# Learning to grok: Emergence of in-context learning and skill composition in modular arithmetic tasks

### v1

@@ 5.2 Attention Heads Implement Essential Skills
We further highlight the importance of the signatures we find in the above paragraphs via comparison with a model that does not generalize o.o.d, in Figure 6(c, d). Note that as the number of pre-training tasks $n_{\mathrm{i.d.}}$ decreases, the attention map similar to the one in panel (a) starts to be mosaicked; and at the same time the PCA projections lose their shape. As a result, those models also lose the ability to perform ICL on modular arithmetic out-of-distribution.
@@
From Figure 5, we infer that the $d=2$ model has only partially acquired skill III, due to its limited capacity. On the other hand, $d=4$ model is very good at combining equations via skill III, explaining its superior performance.
@@ 6 Discussion
We have investigated the emergence of ICL learning and skill composition in AMs on a novel algorithmic dataset. The dataset includes a large discrete set of modular arithmetic tasks and was specifically designed to force models to learn how to solve a variety of tasks. It consists of learning linear modular functions, where the model is expected to identify and perform a modular operation in-context. When the number of tasks is relatively small, the models simply memorize the task vectors and classify the input vectors by the task vectors they have memorized. In this case, the models still have to learn how to perform the memorized tasks because they do generalize in-distribution and develop ICL capabilities. Once the number of training tasks becomes too large the models transition to a qualitatively different, algorithmic approach, where the task vector is determined at inference time.
Finally, we have investigated the learnt representations and showed that qualitatively different circuits are formed in different phases.
@@
We have limited ourselves to algorithmic datasets. Without further investigation, it is difficult to say what lessons can be transported to the realistic language models, and what lessons are specific to the current setting. Mechanistic interpretability analysis of deeper models proved to be much more difficult than that of a smaller models. Consequently, we still do not understand the role of every part of the network in the deeper cases.
@@ Acknowledgments
T.H. thanks Yue Xu and Dayal Singh Kalra for helpful discussions. A.G.’s work at the University of Maryland was supported in part by NSF CAREER Award DMR-2045181, Sloan Foundation and the Laboratory for Physical Sciences through the Condensed Matter Theory Center. The authors acknowledge the University of Maryland supercomputing resources (http://hpcc.umd.edu) made available for conducting the research reported in this paper.
@@ Appendix E Additional Interpretability Results
In this section, we show additional results on interpretability. Where we first show evidence in LABEL:subsecapp:pattern that a $d=4$ model also implements a version of the scale-and-combine algorithm we proposed in Section 5, and it can do better than the $d=2$ model. Then, we continue our discussion on how the algorithm might be implemented inside the model. Importantly, This includes the role of embedding (Section E.1), the role of other attention heads (Section E.2), and finally, the role of MLP and LayerNorm (Section E.3).
@@ E.1 PCA over Embeddings
First we focus the embedding of $d=2$ models. As shown in Figure 9, clearly, the logarithm of each number is split into even and odd groups, and each group forms one clock, which is a suitable embedding for doing modular multiplications. However, one should note that this will not obviate the importance of the head shown in Figure 6(a), as the model still needs a way to distinguish the same number that appears in the different positions in the context.
Curiously, we could not find such a structured embedding for the $d=4$ model, as shown in Figure 10. However, as we will show in the next subsection, this non-structural embedding, together with the first layer, prepares a foundation for the essential heads in the latter layer to create a similar “clock-of-clock" feature as we shown in Figure 6 (a, b).
@@ E.2.1 $d=4$ Model
To continue the story, we analyse the attention heads in different models. We first study the $d=4$ model, where we also find a similar head with “clock-of-clocks" (Figure 6(a)). Moreover, two other heads put together are seemingly equivalent to Figure 6(b). We surmise that any model needs that solves modular arithmetic in context requires such heads.
From Figure 11(a), we see that the head still pays attention locally within three token positions. Importantly, it also creates a clock-of-clock while performing PCA over concatenated $(x,y)$ features. The difference here is that the clock winds twice to go back to its origin. We believe that this factor of two differences in the period means that this head is effectively combing the row of the embedding layer of the $d=2$ model and the role of the head in Figure 6(a). Overall, we do not yet know why the model needs to split the logarithm of numbers into even and odd groups.
From Figure 11(b, c), we find two heads that are structured but different from Figure 6(b). The PCA pattern forms clocks while annotated within logarithm space, but depends on the choice of the sequences. This hints that those two heads are trying to re-scale the examples by comparing them with the previous inputs. We think that the two can be combined to form a similar pattern as Figure 6(b).
@@ E.2.2 $d=2$ Model
Next we focus on the $d=2$ model. In Figure 12, we plot similar PCA to the one in Figure 6(a), but with different concatenated features from the same heads: $(x,z)$ and $(y,z)$. We again see clocks-of-clock, solidifying our argument that this head provides proper spaces for later heads to perform modular operations.
@@
Finally, in Figure 13, we dump all the attention patterns we had for $d=2$ models. The corresopnding PCA over concatenated $(x,y)$ features are shown in Figure 14 (except for layer 1 head 4, for which we plot for $(x,z)$). The special head we choose here has a similar behavior to the one in Figure 6(a), with its main focus shifted by one position but still within two preceding tokens. We believe this head implements another variant of Equation 2 – it re-scales equations for the next layer, then helps them figure out how to combine them to predict new examples.
We have reasons to believe that the non-clock heads are not essential to the models’ performance. However, we leave a careful exploration of this front for future work.
@@
E.3 MLP and LayerNorm
The arguments in Section 5 and the experiments in the previous subsection, strongly suggest that skill III is implemented within the MLP layer. We tried to hunt for signals similar to those in Nanda et al. [19], Gromov [11] from MLP layers, but did not discover conclusive evidence. Similarly, no obvious signals were found in LayerNorm. We think it is crucial to study these layers more carefully. We leave this task for future work.
E.4 Label Noise

### v2

@@ 5.1 Attention Heads Implement Skill I
We further highlight the importance of the structure we find in the above paragraphs via comparison with models, trained with lower pre-training task diversity, that do not generalize o.o.d (Figure 6 c, d). Note that as the number of pre-training tasks $n_{\mathrm{i.d.}}$ decreases, the attention map starts to get mosaicked (top panels); and the PCA projections lose their shape (bottom panels). As a result, these models lose the ability to perform ICL on modular arithmetic out-of-distribution.
5.2 MLPs Implement Skill II and III
After applying Skill I, the model maps numbers onto circular feature spaces, enabling discrete modular operations such as multiplication, division, addition, and subtraction through subsequent model components. To analyze this behavior, we compute the cosine-similarity between layer $l$ (i.e. $l^{th}$ Transformer block) $k$-shot output vectors (standardized) at token position $y$: $\bm{h}_{y}^{l}(x_{1},y_{1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{1}},\cdots,x_{k},y_{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}},x,y)$.
$\cos(\bm{h}^{l}_{y}(x,y),\bm{h}^{l}_{y}(x^{\prime},y^{\prime}))=\frac{\bm{h}^{l}_{y}(\dots,x_{k},y_{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}},x,y)\cdot\bm{h}^{l}_{y}(\dots,x_{k},y_{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z^{\prime}_{k}},x^{\prime},y^{\prime})}{\|\bm{h}^{l}_{y}(\dots,x_{k},y_{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}},x,y)\|\|\bm{h}^{l}_{y}(\dots,x_{k},y_{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}},x^{\prime},y^{\prime})\|}$ (2)
Figure 8: MLP activations (post-ReLU) wrt inputs $(x,y)$
We evaluate this similarity metric across all possible input pairs $(x,y)$ and $(x^{\prime},y^{\prime})$, resulting in a $p^{2}\times p^{2}$ matrix. We also repeat the analysis to layer output vectors at $z$ tokens $\bm{h}^{l}_{z}(\cdot)$. In line with our previous analysis, we use controlled (fixed) in-context examples $(x_{1}\;y_{1}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{1}}\;\dots\;x_{k}\;y_{k}\;{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}z_{k}})$.
For the $d=4$ model, the top panel of Figure 7 reveals a distinctive kaleidoscopic pattern [^3] Patterns are cropped due to file size limit, check the GitHub repo for the full kaleidoscopic pattern. of high cosine-similarities. The pattern corresponds to input pairs $(x,y)$ that share the same ratio $y/x\;\mathrm{mod}\;p$. In the MLP modules, we find highly structured activations as functions of inputs $(x,y)$ (Figure 8 left). Together with our earlier analysis of Skill I, these observations suggest that the MLP layer following the attention module with a “circle-of-circles” head (layer 2, head 4 for $d=4$ model, see Figure 13 (a)) implements division operations over $\mathrm{GF}(p)$. The bottom panel of Figure 7 succinctly demonstrates the algorithmic shift from Ratio Matching to Modular Regression as the number of in-context examples increases. At 2-shot, layer 4 features collapse to identical vectors except when the ratio $y/x$ matches one of the ratios ($y_{1}/x_{1}$ or $y_{2}/x_{2}$). At higher shots, we see a transition to a task-dependent pattern where features align for input pairs $(x,y)$ for which $a\,x+b\,y\;\mathrm{mod}\;p$ match – a signature of Modular Regression.
For $d=2$ models, we again observe the kaleidoscopic pattern in cosine-similarities (Figure 7 top-right) and structured MLP activations (Figure 8 right). However, in this case we only find signatures of Ratio Matching in the cosine similarities (Figure 7 bottom-right), as expected.
@@ 5.2 MLPs Implement Skill II and III
Thus, we conclude that the $d=2$ model has scarcely acquired skill III, due to its limited capacity. On the other hand, $d=4$ model is very good at combining equations via skill III, explaining its superior performance. For a more detailed discussion, see Appendix E.
@@ 6 Discussion
We have investigated the emergence of in-context learning and skill composition in autoregressive models on a novel algorithmic dataset. The dataset includes a large discrete set of modular arithmetic tasks and is specifically designed to force models to learn how to solve a variety of tasks. It consists of learning linear modular functions, where the model is expected to identify and perform a modular operation in-context. When the number of tasks is relatively small the models can only generalize in-distribution. Although such models develop ICL capabilities, they simply memorize these task vectors and use them to classify the input vectors. Once the number of training tasks becomes too large, the models transition to a qualitatively different algorithmic approach, where the task vector is determined at inference time.
Finally, we have examined the learnt representations and shown that qualitatively different circuits are formed in different phases. In the o.o.d. generalization phase, we explain the learnt algorithms and highlight an in-context algorithmic shift in deeper models.
@@
We have limited this work to algorithmic datasets. It remains to be investigated what lessons can be translated to realistic language models, and what lessons are specific to the current setting. White-box interpretability analysis of deeper models has proved to be much more difficult than that of shallower models. Consequently, we still do not understand the role of every individual component of the network in the deeper cases.
@@ Acknowledgments
T.H. thanks Yue Xu and Dayal Singh Kalra for helpful discussions. A.G.’s work at the University of Maryland was supported in part by NSF CAREER Award DMR-2045181, Sloan Foundation. The authors acknowledge the University of Maryland supercomputing resources (http://hpcc.umd.edu) made available for conducting the research reported in this paper.
@@ Appendix E Additional Interpretability Results
In this section, we show additional results on interpretability. Where we continue our discussion on how the algorithm might be implemented inside the model. Importantly, This includes PCA analysis of embedding (Section E.1), other attention heads (Section E.2), attentions and MLPs outputs (Section E.3), and finally, the role of MLPs (Section E.4). We think it is also crucial to study LayerNorm layers, which we leave for future work.
@@ E.1 PCA over Embeddings
First we focus the embedding of $d=2$ models. As shown in Figure 11, clearly, the logarithm of each number is split into even and odd groups, and each group forms one circle, which is a suitable embedding for doing modular multiplications. However, one should note that this will not obviate the importance of the head shown in Figure 6(a), as the model still needs a way to distinguish the same number that appears in the different positions in the context.
Curiously, we could not find such a structured embedding for the $d=4$ model, as shown in Figure 12. However, as we will show in the next subsection, this non-structural embedding, together with the first layer, prepares a foundation for the essential heads in the latter layer to create a similar “circle-of-circles” feature as we shown in Figure 6 (a, b).
@@ E.2.1 $d=4$ Model
To continue the story, we analyse the attention heads in different models. We first study the $d=4$ model, where we also find a similar head with “circle-of-circles” (Figure 6(a)). Moreover, two other heads put together are seemingly equivalent to Figure 6(b). We surmise that any model that solves modular arithmetic in-context requires such heads.
From Figure 13(a), we see that the head still pays attention locally within three token positions. Importantly, it also creates a circle-of-circles while performing PCA over concatenated $(x,y)$ features. The difference here is that the circle winds twice to go back to its origin. We believe that this factor of two differences in the period means that this head is effectively combing the row of the embedding layer of the $d=2$ model and the role of the head in Figure 6(a). Overall, we do not yet know why the model needs to split the logarithm of numbers into even and odd groups.
From Figure 13(b, c), we find two heads that are structured but different from Figure 6(b). The PCA pattern forms circles while annotated within logarithm space, but depends on the choice of the sequences. This hints that those two heads are trying to exchange information with the previous inputs. We leave PCA analysis across different in-context examples to future work.
@@ E.2.2 $d=2$ Model
Next we focus on the $d=2$ model. In Figure 14, we plot similar PCA to the one in Figure 6(a), but with different concatenated features from the same heads: $(x,z)$ and $(y,z)$. We again see circle-of-circles, solidifying our argument that this head provides proper spaces for later heads to perform modular operations.
@@
In Figure 15, we dump all the attention patterns we had for $d=2$ models. The corresopnding PCA over concatenated $(x,y)$ features are shown in Figure 16 (except for layer 1 head 4, for which we plot for $(x,z)$). The special head we choose here has a similar behavior to the one in Figure 6(a), with its main focus shifted by one position but still within two preceding tokens.
In Figure 17, we extend our PCA analysis to different task vectors and up to top-$4$ components, where the spatial structure of the features is better depicted.
We have reasons to believe that the non-circle heads are not essential to the models’ performance. However, we leave a careful exploration of this front for future work.
@@
Figure 17: More PCA analysis similar to Figure 6. Top-$2$ components of PCA contribute $\sim 40\%$ variance of the layer $1$ head and $\sim 20\%$ variance of the layer $2$ head, while the top-$4$ components contribute $\sim 60\%$ and $\sim 40\%$ accordingly.
E.3 PCA over Attention and MLP Outputs
In Figure 18, we plot $d=2$ PCA analysis up to top-$4$ components for attention and MLP outputs.
Figure 18: PCA analysis of Attention and MLP outputs. Attention modules contribute $\sim 50\%$ and $\sim 30\%$ to top-$4$ components of PCA of layer $1$ and layer $2$. MLP modules contribute $\sim 10\%$ and $\sim 12\%$ accordingly.
E.4 Cosine-similarity analysis for MLPs
As we mentioned in Section 5.2, Skills II and III are likely implemented within the MLP layer. While we provided extensive details on how the signature of Skill II was observed in MLP, we could not find clear signals similar to those in Nanda et al. [20], Gromov [11] from MLP layers on Skill III.
Figure 19: MLP (hidden) activations as a function of inputs $(x,y)$ for $d=2,4$ models.
E.4.1 $d=4$ model
Here, we show an extended version of Figure 7 for all layers from the $d=4$ model. We plot for 2-shot (Figure 20) and 16-shot (Figure 21), the task vector used is $(a,b)=(2,6)$.
(a) $y$, layer 1
(b) $y$, layer 2
(c) $y$, layer 3
(d) $y$, layer 4
(e) $z$, layer 1
(f) $z$, layer 2
(g) $z$, layer 3
(h) $z$, layer 4
Figure 20: Cosine-similarity of $d=4$ model for every block, measured at 2-shot.
(a) $y$, layer 1
(b) $y$, layer 2
(c) $y$, layer 3
(d) $y$, layer 4
(e) $z$, layer 1
(f) $z$, layer 2
(g) $z$, layer 3
(h) $z$, layer 4
Figure 21: Cosine-similarity of $d=4$ model for every block, measured at 16-shot.
E.4.2 $d=2$ model
Here, we show an extended version of Figure 7 for all layers from the $d=2$ model. We plot for 2-shot (Figure 22) and 10-shot (Figure 23), the task vector used is $(a,b)=(2,6)$.
(a) $y$, layer 1
(b) $y$, layer 2
(c) $z$, layer 1
(d) $z$, layer 2
Figure 22: Cosine-similarity of $d=2$ model for every block, measured at 2-shot.
(a) $y$, layer 1
(b) $y$, layer 2
(c) $z$, layer 1
(d) $z$, layer 2
Figure 23: Cosine-similarity of $d=2$ model for every block, measured at 10-shot.
E.5 Label Noise